(xy2+ x -2y +3) dx ( x2y -2x -2y) dy =0


1
Expert's answer
2022-02-01T10:02:04-0500

Let us solve the differential equation

"(xy^2+ x -2y +3) dx +( x^2y -2x -2y) dy =0"

Taking into account that "\frac{\partial (xy^2+ x -2y +3)}{\partial y}=2xy-2=\frac{\partial (x^2y -2x -2y)}{\partial x},"

we conclude that the equation is exact, and consequently there exists the function "u=u(x,y)" such that

"\frac{\partial u}{\partial x}=xy^2+ x -2y +3,\ \frac{\partial u}{\partial y}=x^2y -2x -2y."

Therefore, "u=\frac{1}2x^2y^2+\frac{1}2x^2-2xy+3x+c(y)," and hence

"\frac{\partial u}{\partial y}=x^2y -2x +c'(y)=x^2y -2x -2y."

It follows that "c'(y)=-2y," and thus "c(y)=-y^2+C."

We conclude that thegeneral solution of the differential equation

"(xy^2+ x -2y +3) dx +( x^2y -2x -2y) dy =0" is

"\frac{1}2x^2y^2+\frac{1}2x^2-2xy+3x-y^2=C."